### v1

# Understanding LLM Embeddings for Regression

Abstract
With the rise of large language models (LLMs) for flexibly processing information as strings, a natural application is regression, specifically by preprocessing string representations into LLM embeddings as downstream features for metric prediction. In this paper, we provide one of the first comprehensive investigations into embedding-based regression and demonstrate that LLM embeddings as features can be better for high-dimensional regression tasks than using traditional feature engineering. This regression performance can be explained in part due to LLM embeddings over numeric data inherently preserving Lipschitz continuity over the feature space. Furthermore, we quantify the contribution of different model effects, most notably model size and language understanding, which we find surprisingly do not always improve regression performance.
1 Introduction and Related Work
Regression is a fundamental statistical tool used to model the relationship between a metric and a selected set of features, playing a crucial role in various fields, enabling predictions, forecasting, and the understanding of underlying relationships within data. Traditional regression techniques often rely on handcrafted features or domain-specific knowledge to represent input data. However, the advent of Large Language Models (LLMs) and their ability to instead process semantic representations of text has raised the question of whether regression can instead be performed over free-form text.
Figure 1: Rugged surface of a 5D Sphere function when inputs are represented as Gemini embeddings of dimension 6K+, post-processed by t-SNE into 2D space.
Previous works have predominantly examined the topic of LLM-based regression through decoding, i.e. generating floating point predictions using token-based sampling. For example, (Song et al., 2024) examines the case when the model is fully accessible and fine-tunable against data, while (Vacareanu et al., 2024) study the ability of service-based closed-source LLMs such as GPT-4 using in-context learning.
One understudied case however is the use of service-based LLM embeddings - fixed vector representations derived from pre-trained (but frozen) language models, which are ubiquitously offered among most LLM services (OpenAI, 2023; Google, 2024). Although they are used frequently in recent applications such as retrieval (Karpukhin et al., 2020), semantic similarity (Li et al., 2020), and a variety of other downstream language tasks (Liu et al., 2020), there has been very little research for their general and direct use in regression, outside of specific applications such as Bayesian Optimization (Nguyen et al., 2024).
In contrast to decoding-based regression techniques, embedding-based regression allows the possibility of cheap data-driven training using inexpensive and customizable post-embedding layers such as multi-layer perceptrons (MLPs). However, as shown in Figure 1, when the domain of a simple function is expressed using high-dimensional embeddings, unexpected characteristics and irregularities can arise, prompting the need for a thorough analysis. Furthermore, LLMs by default are not explicitly trained for embedding-based regression, rather purely for token generation, and thus it is worth analyzing the emergent behaviors of LLMs when applied to embedding-based regression.
This paper investigates the behavior of these LLM embeddings when used as features for traditional regression tasks. Most notably, our findings are:
  • LLM embeddings are dimensionally robust, i.e. regression performance can remain strong even over high-dimensional data, whereas traditional representations significantly suffer.
  • Over numeric formats, LLM embeddings preserve Lipschitz-continuity and smoothness over feature space, which naturally enables regression when using a downstream MLP head.
  • Factors which directly impact language understanding (e.g. size, pre-training, and input formatting) have more nuanced effects for regression and do not always provide significantly better outcomes.
2 Problem and Methodology
A regression task $\mathcal{T}=(f,\mathcal{X},\mathcal{D})$ consists of an underlying scalar-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ over an input space $\mathcal{X}$. Provided are offline training data $\mathcal{D}_{train}=\{(x_{1},y_{1}),...,(x_{T},y_{T})\}$ collected from querying $f$ and an analogous test set $\mathcal{D}_{test}$ for evaluation. Given access to training data $\mathcal{D}_{train}$, the goal is to obtain accurate predictions over test points $(x,y)\in\mathcal{D}_{test}$, usually measured by an aggregate performance measure, e.g. mean squared error or Kendall-Tau ranking scores.
Required by nearly all learnable regression methods are features, which we assume come from an embedder $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$ which takes an input $x$ and returns a fixed-dimensional feature representation, of dimension $d$. Here, we use the terms "features" and "embedding" interchangeably, since traditional methods typically use a canonical, manually defined feature engineering method for tabular data, in which continuous values are normalized and categorical selections are one-hot encoded. This feature vector $\phi(x)$ is then sent to a downstream predictor, e.g. MLP or random forest, which is trained using a loss function such as mean squared error.
Language models also provide a canonical definition of embedding, which typically consists of, in order:
  1. 1.
    Tokenizing a string representation $x$ into $L$ tokens.
  2. 2.
    Obtaining a "soft prompt" $\mathbb{R}^{L\times v}$ via vocabulary look-up.
  3. 3.
    Applying a forward pass of a Transformer to obtain an output $\mathbb{R}^{L\times f}$.
  4. 4.
    Pooling down to a fixed dimension vector in $\mathbb{R}^{d}$.
Afterwards, one may also attach an MLP predictor head and apply an analogous training procedure as in the traditional case. Thus we can see that the only difference becomes the input representation $\phi$, i.e. whether we used a traditional $\phi_{\text{trad}}$ or LLM-based $\phi_{\text{LLM}}$.
While it is straightforward to assume that the whole process outlined for LLMs should constitute the definition of a language model embedding $\phi_{\text{LLM}}$, it is not obvious how much each of these steps may contribute to the final regression result. For instance, one could simply skip applying a forward pass in step (3) and pool the soft prompt directly, or use a randomly initialized model as opposed to a pretrained one. We extensively study this case in Section 3.3.
2.1 Modeling
To minimize confounding factors and maintain fairness during comparisons, we use the exact same MLP prediction head (2 hidden layers, ReLU activation), loss (mean squared error), and $y$-normalization scheme (shifting by empirical mean and dividing by empirical deviation), regardless of using $\phi_{\text{LLM}}$ and $\phi_{\text{trad}}$. Note however, that the embedding dimensions of the two representations may be different, and so we distinguish them using notation $d_{\text{llm}}$ and $d_{\text{trad}}$ respectively, where typically $d_{\text{llm}}\gg d_{\text{trad}}$. Further details can be found in Appendix B.1.
To demonstrate consistent results over different families of language models, we benchmark over both the T5 (Raffel et al., 2020) and Gemini 1.0 (Google, 2024) families, which use different architectures (encoder-decoder and decoder-only), different vocabulary sizes (32K and 256K), and embedding dimensions (See Appendix B.2) respectively. However, to remain consistent with the definition of embedding, we follow previous literature (Li et al., 2020; Reimers and Gurevych, 2019) and use average-pooling as the canonical method of aggregating Transformer outputs, and thus the embedding dimension $d_{\text{llm}}$ is equivalent to the the output feature dimension $f$ following a forward pass.
Similar to previous work (Song et al., 2024; Nguyen et al., 2024), for string representations of $x$ from any regression task, by default we use a key-value JSON format with consistent ordering of keys, i.e. {param1:value1,param2:value2,…}, with specific examples shown in Appendix C.
2.2 Regression Tasks
For regression tasks, we first use synthetic, closed-form objective functions in order to produce controlled studies in which we may query any $x$ from the input space. Our synthetic functions are defined from the standard Black-Box Optimization Benchmarking (BBOB) suite (Elhara et al., 2019). To avoid confounding terminology between embedding "dimension" $d$ and the intrinsic "dimension" of an objective $f$, we denote the latter as "degree-of-freedom" (DOF), and thus $f(\cdot)$ is dependent on input coordinates $x^{(1)},\ldots,x^{(\textrm{DOF})}$, each of which is between $[-5,5]$. This provides a comprehensive variety of both convex and non-convex objective landscapes to regress upon.
We further use real-world regression tasks representative of those encountered in the wild and in industry settings by benchmarking over offline objective evaluations found in Google Vizier (Golovin et al., 2017), which optimizes Google’s largest production and research systems. These consist of four families, with each family containing at least 50 individual yet similar regression tasks. The families are:
  • AutoML (Google Cloud, 2023): Automated Machine Learning platform for automating TFX (Google, 2023) pipelines (e.g. batch size, activation, layer counts) over tabular or text data.
  • Init2Winit (Dahl et al., 2023): Learning rate scheduling parameters influencing common image classification tasks (e.g. ResNets on CIFAR-10 and ImageNet).
  • XLA (Phothilimthana et al., 2021): Tuning for the Accelerated Linear Algebra (XLA) compiler which affects LLM serving latencies.
  • L2DA (Yazdanbakhsh et al., 2021): "Learning to Design Accelerators", for improving accelerators such as TPUs and corresponding computer architectures to improve hardware performance.
In the real world regression tasks, each parameter may be continuous or categorical, and we define the DOF of such a task by its number of parameters. Note that for synthetic objectives, where all inputs are continuous, $d_{\text{trad}}=\textrm{DOF}$. However, for real-world tasks with categorical parameters, $d_{\text{trad}}>\textrm{DOF}$ due to additional one-hot encodings.
For obtaining data, we may either sample $(x,y)$ pairs (in the case of synthetic objectives where $x$ are uniformly sampled from $\mathcal{X}$), or use the given offline data (in the case of real-world tasks, where they were actual evaluations from an optimization trajectory), using a standard 8-1-1 train-validation-test split.
Due to the inherent differing of metric scales across tasks, it would be inappropriate to aggregate results based on scale-dependent metrics such as mean squared error (MSE). Furthermore, we found that the selection of the regression metric (e.g. Kendall-Tau, Pearson, mean squared error, mean absolute error) did not matter for comparisons, as they all strongly correlated with each other. Thus, by default we report the Kendall-Tau ranking correlation, which is always within $[0,1]$ and can also be aggregated across different tasks.
3 Experimental Results
3.1 High Dimensional Regression
We begin by demonstrating cases in which LLM embeddings better represent inputs over high degree-of-freedom spaces than traditional representations. In Figure 2, we show that for a subset of functions, LLM embeddings possess surprising robustness, retaining the same performance for varying DOFs whereas traditional baselines such as XGBoost and MLPs significantly falter over higher DOFs.
Figure 2: Higher ($\uparrow$) is better. Degrees of freedom (DOF) vs Kendall-Tau correlation for various BBOB functions. Results are averaged over 12 runs for each regression method. Each task’s data consists of 500 $(x,y)$ evaluations sampled uniformly across the search space, using a 8-1-1 split for train-validation-test.
This result is not universal however, as we show in Appendix A.1, this pattern does not apply for a few selected functions, but nonetheless it occurs in the majority of the BBOB functions. We further corroborate this observation over real-world tasks in Table 1. We see that in general, regressions on LLM embeddings outperform traditional methods more often for tasks with higher DOFs (AutoML and XLA).
Task Name Avg. DOF T5-Small % T5-XXL % Gemini Nano % Gemini Pro %
Init2Winit 4 6.7 8.0 11.3 19.0
L2DA 10 2.7 12.0 9.3 10.7
AutoML 29 30.7 41.3 29.3 36.0
XLA 35 17.2 29.3 18.9 24.1
Table 1: Percentage of tasks in which $\phi_{\text{LLM}}$ outperforms $\phi_{\text{trad}}$ across various real world regression tasks. Results reported for 75 tasks per family, except for XLA, which only contains 58 tasks. Full results in Appendix A.2.
3.2 LLM Embedding Smoothness
Particularly due to the discrete nature of tokenization, it is non-obvious whether LLM embeddings possess a notion of continuity in embedding space. For example, assuming character-wise tokenization, 1.234 is not so numerically distant from 1.567, but is token-wise distant, as the majority of the tokens (234 and 567) are not shared.
The notion of continuity and smoothness is crucial for neural network generalization (Kalimeris et al., 2019; Neyshabur et al., 2018), robustness (Weng et al., 2018), vulnerability to adversarial examples (Goodfellow et al., 2015), and more. We can characterize smoothness in the regression case by the Lipschitz-continuity induced by a representation $\phi$ in its latent space $\mathbb{R}^{d}$.
Figure 3: Left-skewness ($\leftarrow$) is better. NLFDs induced by $\phi_{\text{LLM}}$ (Gemini Nano) and $\phi_{\text{trad}}$. Top: Cases where $\phi_{\text{LLM}}$ outperforms $\phi_{\text{trad}}$ for regression. Bottom: Vice-versa where $\phi_{\text{trad}}$ outperforms $\phi_{\text{LLM}}$.
[…]
In order to analyze the continuity of an embedding $\phi$ with respect to offline data $\mathcal{D}$, we define a Normalized Lipschitz Factor Distribution (NLFD) as follows:
  1. 1.
    Normalize each embedding vector $\phi(x)$ coordinate-wise to have zero mean and unit variance across the dataset $\mathcal{D}$.
  2. 2.
    For each $x\in\mathcal{D}$, choose $x^{\prime}\in\mathcal{D}$ such that $\phi(x^{\prime})$ is the nearest $\ell_{2}$ neighbor of $\phi(x)$, and compute the Lipschitz factor $L(x,x^{\prime})$.
  3. 3.
    To assume an average embedding norm of $1$ for different embedding dimensions $d$, we downscale all Lipschitz factors by $\sqrt{d}$.
We see that there is a high inverse relationship between the skewedness of the NLFD and regression performance. Specifically, in Figure 3, when $\phi_{\text{LLM}}$ outperforms $\phi_{\text{trad}}$ for regression, $\phi_{\text{LLM}}$’s distribution of Lipschitz factors also tends to skew relatively more to zero than $\phi_{\text{trad}}$, and vice-versa.
Figure 4: Relationship between gaps in NLFD (via Z-score) and regression performance for all 23 BBOB functions. Relationship is quantified using (K, S, P), which respectively are Kendall-Tau, Spearman and Pearson correlations. Top: We vary model size within the T5 model family. Bottom: We vary the objective’s DOF for Gemini Pro.
To formally quantify comparisons between NLFDs from $\phi_{\text{LLM}}$ and $\phi_{\text{trad}}$, for a fixed regression task, we may thus compute the Z-score using the difference of the two distributions:
$Z=\frac{\mu_{\phi_{\text{trad}}}-\mu_{\phi_{\text{LLM}}}}{\sqrt{\sigma_{\phi_{\text{trad}}}^{2}+\sigma_{\phi_{\text{LLM}}}^{2}}}$ (1)
where $\mu_{\phi}$ and $\sigma_{\phi}$ are respectively mean and standard deviations of the NLFD of a representation $\phi$. We may then observe the relationship between gaps in representation smoothness vs. regression performance. In Figure 4 with extended results in Appendix A.3, we see that for a given BBOB regression task, the Z-score (i.e. gap in embedding smoothness) is highly correlated with the gap in regression performance, regardless of the model used (T5 or Gemini) or the DOF of the underlying objective $f$.
We further visualize whether $\phi_{\text{LLM}}$ is distance aware, i.e. whether $\phi_{\text{LLM}}(x)$ are $\phi_{\text{LLM}}(x^{\prime})$ are close in embedding space if $\phi_{\text{trad}}(x)$ and $\phi_{\text{trad}}(x^{\prime})$ are close. As mentioned before however, there is no ground truth notion of "closeness" - nonetheless, we use $\phi_{\text{trad}}$ as a point of comparison. Since it is inappropriate to simply sample $x$’s uniformly in a high DOF space, as then average distances concentrate around $\sqrt{\textrm{DOF}}$, we instead take a reference point and sample points from $\ell_{2}$-balls of increasing distance from the reference.
In Figure 5, we see that distances over the LLM embedding space are correlated with the traditional measure of distance, but may be non-linearly warped, which benefits LLM-based regression in certain cases as seen in Section 3.1.
Figure 5: t-SNE for Gemini (Nano and Pro) embeddings of points sampled around a DOF=100 reference point. Traditional $\ell_{2}$ distance is overlayed in color.
3.3 Model Effects
In this subsection, we comprehensively investigate the impact of many common LLM factors on regression performance.
Are Larger Models Always Better? Within the research community, the prevailing assumption is that there exists a direct correlation between language model size and performance improvement. However, with the rise of leaderboards such as LMSYS (LMS, 2023), smaller models have been shown to outperform larger competitors, due to differences in their "recipe", such as training data quality, pre-training and post-training techniques, and architecture.
Figure 6: Higher ($\uparrow$) is better. Model size vs regression performance on hyperparameter tuning tasks across T5 and Gemini model families. Median performance is plotted, along with 40-60 percentiles as error bars.
In Figure 6, we see that over various real world regression tasks, T5 models exhibit a clear trend of improved performance when increasing model size, when training methodology is fixed. In contrast, model tiers within the Gemini family exhibit substantial variance, and larger model sizes do not consistently translate to superior results. We hypothesize this is due to differences in Gemini "recipes", as e.g. different model tiers may have used different pre-training datasets, architecture tweaks, and post-training configurations, whereas all T5 model sizes have only been pre-trained on the C4 web crawl corpus.
Does Language Understanding Actually Help? Recent works (Li et al., 2020; Devlin et al., 2019) have claimed that logit-based embeddings mostly measure the semantic similarity between string inputs, and thus it is unconfirmed whether they may be beneficial for numeric regression tasks. To resolve this, using the T5 family, we compare against using (1) a randomly initialized model for the forward pass, and (2) representing our features via vocabulary embeddings without applying a forward pass.
Figure 7: Kendall-Tau regression comparisons when comparing to random initialization (left) and vocabulary embeddings (right). Each bar is averaged across 75 tasks per family.
In Figure 7, we see that the default mode of applying a forward pass of a pre-trained model performs the best, as expected. However, it is worth noting that in some tasks such as AutoML and L2DA, the improvement is surprisingly quite minimal, suggesting that applying forward passes by pretrained models does not always help for regression.
Figure 8: Difference in Kendall correlation when using full dictionary containing feature names, or only values.
We further ablate differences in string representation, i.e. whether by default to show feature names as {param1:value1, param2:value2,…} or omit them, only showing [value1,value2,…]. In Figure 8, for the majority of tasks, omitting feature names does not significantly affect performance, although specific tasks such as XLA do benefit from feature names. This is surprising, as presumably feature names in XLA tasks such as auto_cross_replica_sharding are not as common as names such as batch_size or learning_rate found in both AutoML and Init2winit.
The results of Figures 7 and 8 combined lead to additionally surprising conclusions, such as language-to-numeric transfer. For instance, inputs $x$ from Init2Winit tasks only possess numeric values, and as expected, removing feature names does not significantly change regression results. Yet applying forward passes by pre-trained T5 models still benefits regression, despite the fact that T5’s pre-training data contains mostly web-corpus data which is unlikely to contain significant amounts of scientific or numeric information (Dodge et al., 2021).
More Training Data Reduces Baseline Gaps: Intuitively, as more samples are available in a task, the difference in inductive biases between regression methods should matter less, since predictions will be more influenced by training data. We verify this in Figure 9, where we see that for tasks with low numbers of $(x,y)$ points, there is more variability in performance between using $\phi_{\text{LLM}}$ and $\phi_{\text{trad}}$, but additional training points decreases these differences.
Figure 9: Performance gap between an MLP baseline and regression over T5-XXL embeddings for individual trials within the AutoML and XLA task settings. Higher ($\uparrow$) is better for LLM embeddings. Error bars are plotted for $\{0.5,1.0,2.0\}$ of the standard deviation.
4 Conclusion and Future Work
We thoroughly investigated multiple important aspects around the use of LLM embeddings for traditional regression. We found that LLM embeddings can be quite performant for input spaces with high degrees of freedom, and proposed the Lipschitz factor distribution to understand the embedding-to-objective landscape and its relationship to regression performance. We further investigated the nuanced conditions for which better language understanding does improve LLM-based regression.
Since strings, and more generally tokens, can represent many varieties of data, it is worth further understanding the effects of LLM embeddings over non-tabular forms of inputs, including combinatorial objects such as graphs, and even other modalities such as images and videos.
Acknowledgments
We thank Yutian Chen, Daniel Golovin, Chansoo Lee, Tung Nguyen, and Sagi Perel for relevant discussions during experimentation and the writing of this paper. We further thank the organizers of the Google DeepMind Academy Program for providing the opportunity to do this research.
Appendix
Appendix A Extended Experiments
A.1 High Dimensional Regression
For full transparency, In Figure 10, we display BBOB functions where LLM-based regression was not consistently dimensionally robust against MLP and XGBoost baselines. Note that even in these cases, we still see certain cases where a language model outperforms at least one of the baselines, e.g. in the Discus and DifferentPowers functions, Gemini and T5 outperform MLP but not XGBoost.
Figure 10: Following Figure 2 in the main body, we present BBOB functions in which LLM embeddings did not completely outperform traditional baselines.
A.2 Real World Results
Figure 11: Full Results over real world tasks. Displayed is the mean Kendall-Tau Correlation over all tasks within each family.
A.3 Performance Correlations
Following Figure 4, in Table 2, we see that the relationship between the smoothness induced by the embedding and the performance in regression is consistent throughout.
Model DOF=5 DOF=10 DOF=25 DOF=50 DOF=100
Gemini Nano 0.81 0.81 0.70 0.75 0.86
Gemini Pro 0.78 0.77 0.72 0.82 0.88
T5-Small 0.75 0.76 0.79 0.79 0.76
T5-Large 0.78 0.73 0.79 0.85 0.79
T5-XL 0.82 0.60 0.80 0.86 0.85
T5-XXL 0.72 0.76 0.82 0.83 0.83
Table 2: Full set of data for Pearson correlation $\rho$ between Kendall’s regression performance and gap in NLFD between input and embedding space for regression on all 23 BBOB functions, over DOF=$[5,10,25,50,100]$.
Appendix B Exact Modeling Details
B.1 Hyperparameters Used
The full list of hyperparameters and training details for MLP-based regression (using traditional and language model features):
  • Regression Head: MLP with 2 ReLU hidden layers of dimension 256.
  • $y$-Normalization: We compute the empirical mean $\mu$ and standard deviation $\sigma$ over all $y$-values in the task’s training data, and apply $y\leftarrow(y-\mu)/\sigma$ as a preprocessing step.
  • Optimizer: AdamW with sweeped learning rates across {1e-4, 5e-4, 1e-3, 5e-3, 1e-2} and weight decay across {0, 1e-1, 1}.
  • Loss: Mean Squared Error.
  • Maximum Epochs: 300, with early stopping enabled.
For XGBoost, we additionally grid-searched over the following parameters for each task:
  • ‘‘min_child_weight": [1, 5, 10]
  • ‘‘learning_rate": [0.001, 0.01, 0.1]
  • ‘‘gamma": [0.0, 0.3, 0.5]
  • ‘‘subsample": [0.6, 0.8, 1.0]
  • ‘‘colsample_bytree": [0.6, 0.8, 1.0]
  • ‘‘max_depth": [3, 5, 7]
B.2 Embedding Sizes
Table 3 displays the embedding $d_{\text{llm}}$ for each model used in our experiments. As mentioned in the main text, note that $d_{\text{llm}}$ is significantly larger than $d_{\text{trad}}$.
T5 Model $d_{\text{llm}}$
Small 512
Large 1024
XL 2048
XXL 4096
Gemini Model $d_{\text{llm}}$
Nano 1536
Pro 6144
Ultra 14336
Table 3: Embedding dimensions $d_{\text{llm}}$ for T5 and Gemini model families.
Appendix C Example String Representations
Below are example string representations of $x$ for different regression task families. ‘…’ denotes that there are actually more parameters, but we omit them due to page limits.
Task Family Example Representations
BBOB x0:0.32, x1:-4.21, x2:3.12, x3:1.56
AutoML batch_size:128, ml_feature_selection_threshold:0.05, model_type:‘DNN_ESTIMATOR’, activation_fn:‘selu’, batch_norm:‘False’, bucketization_strategy:‘mdl’, dropout:0.071, hidden_units:359
Init2Winit lr_hparams.base_lr:0.0696, opt_hparams.0.hps.one_minus_b1:0.2823, opt_hparams.0.hps.one_minus_b2:0.0432, opt_hparams.1.hps.weight_decay:0.0023
XLA auto_cross_replica_sharding:‘False’, rematerialization_percent_shared_memory_limit:97, spmd_threshold_for_windowed_einsum_mib:100000, ...
L2DA input_activation_memory_depth:11.0, instruction_memory_depth:15.0, io_bandwidth_gbps:4.321, narrow_memory_capacity_bytes:21.0, ...